# Nicholas Michel    Mohamed E. Najd    Ghada Almashaqbeh

### v3

@@ 2 Background
We provide an overview of the general functionality of AMMs and the chainBoost framework that we use in the design of $\mathsf{ammBoost}$.
@@
Automated market makers. AMMs build platforms for token trading powered by the users themselves. This is done by establishing liquidity pools such that a pool trades a pair of tokens, say tokens $A$ and $B$. Users are divided into: clients which could be sellers and buyers, and liquidity providers (LPs). Providing liquidity comes from the sellers themselves since buying token $A$ requires paying the price using token $B$ (and vice versa), and from LPs who deposit tokens in the pool and collect fees in return.
Constant function market makers (CFMM) is a popular implementation choice for computing the trading price in AMMs. This formula keeps the ratio of token reserves, and consequently prices, in the pool as balanced as possible to reduce price slippage. In particular, the price of token $A$ multiplied by the price of token $B$ equals a constant. Let the reserves of tokens $A$ and $B$, i.e., their total amounts, in the pool be $res_{A}$ and $res_{B}$, respectively, then the price of token $A$ is $res_{B}/res_{A}$ and the price of token $B$ is $res_{A}/res_{B}$. Accordingly, for an order trading an amount of token $A$, $amt_{A}$, the amount of token $B$, $amt_{B}$, that this order receives is computed as: $amt_{B}=res_{B}-(res_{A}\cdot res_{B})/(res_{A}+amt_{A})$.
@@
At a basic level, an AMM implementation supports several transaction types: for trading, there are (exact input and exact output) swaps, and for liquidity management, there are mints, burns and collects that allow LPs to submit liquidity positions, collect their fees, and withdraw these positions, respectively. AMMs may provide additional services, such as flash loans [69] allowing clients to take advantage of arbitrage opportunities across different platforms. Furthermore, more sophisticated liquidity approaches are being adopted, e.g., concentrated liquidity [44] that enable defining a price range over which liquidity will be applied to address issues related to inefficient use of provided funds.
@@
The functionality of an AMM is commonly implemented as a set of smart contracts on top of a smart contract-enabled blockchain, where Ethereum is the dominant choice so far. These contracts create and manage the liquidity pools, and provide the API needed to interact with the AMM. Residing on a public blockchain led to several financial and security issues [37, 58, 63], e.g., front-running attacks, sandwich attacks, miner/maximal extractable value, etc. Understanding and solving these issues are active research areas. We do not discuss these issues further since they are not the focus of this work; we target the storage cost and throughput of AMMs.
@@
The chainBoost framework. chainBoost [57] is a sidechain-based solution that aims to reduce the blockchain storage footprint and confirmation delays, and boost transaction throughput. It introduces a new sidechain architecture that shares the workload with the mainchain, and enables pruning stale records. As such, this sidechain has a mutual-dependence relation with the mainchain. Transactions are classified into sidechain and mainchain transactions, where all service-related operations that can be summarized go to the sidechain, while the rest stay on the mainchain. The sidechain works in parallel to the mainchain, and operates in epochs and rounds (an epoch is $\omega$ consecutive rounds and a round is the period during which a new block is mined). At the end of each epoch, the mainchain is synced with summaries of the workload processed by the sidechain in that epoch.
Figure 1: The chainBoost framework.
@@
As shown in Figure 1, the sidechain is managed by the mainchain miners, where for each epoch, a committee is elected to process the sidechain traffic during that epoch. The rest of the mainchain miners, who are not on the committee, do not process the sidechain traffic, thus reducing their load. To speedup agreement, chainBoost employs a practical Byzantine fault tolerance (PBFT)-based consensus (similar to those in [52, 48]) for the sidechain. [^5] Similar to chainBoost, to simplify the presentation, we adopt a leader-based PBFT in which a leader proposes a block for the committee to agree on (as in [52]). Nonetheless, voting-based PBFT (as in [48]) can be used instead.
@@
The sidechain is composed of two types of blocks (as shown in the figure): temporary meta-blocks and permanent summary-blocks. For each sidechain round, the committee mines a meta-block containing the transactions they processed, so that once a transaction appears in a meta-block it is considered final. In the last round of the epoch, this committee mines a summary-block summarizing all state changes induced by the meta-blocks within that epoch. After that, it issues a sync-transaction containing the summarized state changes, which the mainchain miners use to update the relevant state variables on the mainchain. Once the sync-transaction is confirmed on the mainchain, all meta-blocks used to produce the respective summary-block are discarded. This significantly reduces the sidechain size, and subsequently, the mainchain size. At the same time, having permanent summary-blocks allows anyone can verify the source of the state changes recorded on the mainchain.
Applicability of dependent-sidechains to AMMs’ setting. Our setting is different from the one in [57]: First, the mainchain and sidechain miners run different protocols. In $\mathsf{ammBoost}$, the mainchain miners belong to a smart contract-enabled blockchain, and the AMM is simply an application deployed on that blockchain. Thus, the sidechain must have its own miner population, to run its protocol, and a technique, such as proof-of-stake, to mitigate Sybil attacks. [^6] Indeed, a miner can choose to operate on both the mainchain and the sidechain. Still this miner runs two protocols, one for each chain, rather than one protocol. Second, in chainBoost the two chains are mutually-dependent, i.e., their security and valid operation depend on each other. In $\mathsf{ammBoost}$, the dependence is unidirectional; interruptions on the mainchain impact the sidechain since the base contract that keeps track of the AMM state resides on the mainchain, but not vice versa. Sidechain interruptions will indeed lead to invalid state of the base AMM contract, but this contract resides on the application layer and does not impact the underlying mainchain or other deployed dApps. Third, in $\mathsf{ammBoost}$, the actual token reside on the mainchain, while the trading and liquidity-related activities are handled by the sidechain. Thus, a mechanism is needed to handle token payouts and deposits to enable accepting and processing only valid transactions. Devising techniques to address these issues resemble the core novelty of $\mathsf{ammBoost}$ system design.
3 Preliminaries
@@
Notation. We use $\lambda$ to denote the security parameter, and $\mathtt{pp}$ to denote the system public parameters. We use $\mathcal{L}$ to denote a ledger (or blockchain), $\mathcal{L}_{\mathsf{mc}}$ to denote the mainchain ledger, and $\mathcal{L}_{\mathsf{sc}}$ to denote the sidechain ledger. The former is the smart contract-enabled blockchain on top of which the AMM base smart contract is deployed, while the latter is the blockchain of the AMM ecosystem. Each party maintains a secret key $\mathsf{sk}$ and a public key $\mathsf{pk}$. Lastly, we use $\mathsf{PPT}$ as a shorthand for probabilistic polynomial time.
@@
System model. $\mathsf{ammBoost}$ involves a base smart contract representing the AMM on the mainchain, and a sidechain that processes most of the AMM workload. Anyone can join/leave the AMM at anytime, and these parties are known using their public keys. Participants are three types: clients $\mathcal{C}$ who are only interested in using the AMM trading services, liquidity providers $\mathcal{LP}$ who provide liquidity for the pools operated by the AMM, and miners $\mathcal{M}$ who maintain the AMM sidechain. We do not place any restrictions on the mainchain beyond being a secure smart contract-enabled blockchain. $\mathsf{ammBoost}$ operates in rounds and epochs (as defined earlier). The sidechain is managed by a committee elected from the sidechain miners, where a new committee is elected for each epoch. This committeee runs a PBFT-based consensus to mine new blocks: temporary meta-blocks that record transactions, and permanent summary-blocks that summarize meta-blocks mined in an epoch. The committee also issues sync-transactions to sync the base AMM smart contract deployed on the mainchain. Accordingly, the $\mathsf{ammBoost}$ framework provides the following functionalities:
@@
Takes as input the the security parameter $\lambda$ and the mainchain $\mathcal{L}_{\mathsf{mc}}$. It configures the system public parameters $\mathtt{pp}$, and deploys a base AMM smart contract on $\mathcal{L}_{\mathsf{mc}}$. It outputs $\mathtt{pp}$ and the initial sidechain ledger state $\mathcal{L}_{\mathsf{sc}}^{0}$ (which is the genesis block referencing the mainchain block containing the base contract).
@@
Takes as input $\mathtt{pp}$ and outputs the initial local state of the party $\mathsf{state}$, which contains a keypair $(\mathsf{sk},\mathsf{pk})$, and in case of miners, the current view of $\mathcal{L}_{\mathsf{sc}}$.
@@
  • $\mathsf{tx}_{\mathsf{collect}}$: Allows an LP to collect fees accrued due to providing liquidity.
  • $\mathsf{tx}_{\mathsf{burn}}$: Allows an LP to withdraw her liquidity.
@@
Takes as input the current sidechain ledger state $\mathcal{L}_{\mathsf{sc}}$, a new block $\mathsf{B}$ with type $\mathsf{btype}=\mathsf{meta}$ or $\mathsf{btype}=\mathsf{summary}$. It outputs 1 if $\mathsf{B}$ is valid based on the syntax/semantics of the block type, and 0 otherwise.
@@
Security model. We aim to develop a secure state growth control solution that preserves the valid and secure operation of the underlying AMM. $\mathsf{ammBoost}$ builds a sidechain, which is basically a blockchain, that interacts with the application layer of the mainchain through the base AMM smart contract. This sidechain must be a secure ledger as defined below.
Safety:
Liveness:
@@
A ledger must record only valid transactions and blocks, thus its protocol is parameterized by predicates to verify transaction and block validity. For dApps, validity is governed by the code of their smart contracts, and miners ensure that the ledger state changes have been produced by a successful execution of this code. $\mathsf{ammBoost}$ reduces the AMM functionality deployed on the mainchain, and it processes most of the workload (following the same logic of the AMM) on the sidechain. Thus, in our security analysis, we show that $\mathsf{ammBoost}$ preserves the security and correct operation (i.e., safety and liveness) of the original AMM.
@@
Adversary model. We assume the mainchain to be secure as defined above. For the sidechain, we have honest miners who follow the protocol, and malicious miners controlled by the adversary who may behave arbitrarily. The adversary can deploy new miners or corrupt existing ones, without going above the threshold of faulty nodes required by the sidechain consensus protocol. The adversary can see all messages and transactions sent in the system (since we deal with public permissionless blockchain systems) and can reorder these messages and delay them. We assume bounded-delay message delivery, so any sent message (or transaction) will be delivered within $\Delta$ time as in [48, 59, 53]. We assume slowly-adaptive adversaries [34] that can corrupt miners only at the beginning of each epoch. Lastly, we deal with $\mathsf{PPT}$ adversaries.
Figure 2: The $\mathsf{ammBoost}$ framework (Txs is an abbreviation for transactions).
4 System Design
@@
$\mathsf{ammBoost}$ changes the AMM deployment structure, as shown in Figure 2. The smart contract on the mainchain is minimal; it mainly tracks the token balances of the liquidity pools and the users, while most of the transaction processing is moved to the sidechain. Summaries of the sidechain processed traffic are used to sync the AMM smart contract on the mainchain. In this section, we present the design of $\mathsf{ammBoost}$ including system setup, architecture and operation, handling interruptions, and its security.
4.1 System Setup
@@
The setup phase, as depicted in Figure 3, mainly specifies the traffic split between the chains and the summary rules for the sidechain traffic, as well as the sidechain parameters such as the epoch length and its consensus configuration (e.g., committee size). [^7] The epoch duration impacts syncing frequency. Short epochs mean more sync-transactions, which incurs more gas cost and may impact throughput—as they are processed by the mainchain, however users would receive their tokens faster compared to long epochs. We empirically study the impact of this parameter in Section 6. Also, this phase involves deploying the AMM base smart contract on the mainchain and creating the sidechain.
Traffic classification and summary rules. AMM transactions and operations are divided into two groups: pool management and trading-related. Creating and managing token pools, as well as dispensing tokens to clients and LPs, are done on the mainchain since these deal with actual tokens. Flashes are also handled by the mainchain since they require instant token dispensing rather than at the end of the epoch (which is the case for any operation processed by the sidechain). The rest of the transactions, including swaps, mints, burns, and collects, are handled by the sidechain.
In $\mathsf{ammBoost}$, the sidechain does not hold custody of tokens, it just tracks their balances based on the processed transactions. Thus, during each epoch, the sidechain produces two structures:
  • A payout list containing users’ public keys and the amount/type of tokens they should receive. This list is simply the updated deposit balance produced at the end of an epoch.
@@
$\mathsf{SystemSetup}(1^{\lambda},\mathcal{L}_{\mathsf{mc}})$: Takes as input the security parameter $\lambda$ and the current mainchain state $\mathcal{L}_{\mathsf{mc}}$, and does the following:
@@ 1. Generate the sidechain configuration parameters:
  • All parameters needed for the sidechain consensus protocol.
@@
2. Deploy the base contract $\mathsf{TokenBank}$ on the mainchain.
@@
The actual token dispensing and deduction happen at the end of an epoch when the sidechain summaries are received. The new state of the pool token balances on the mainchain will be computed based on these lists. Also, the payout encompasses refunding any leftover in the deposits to their owners as will be shown shortly. In Section 4.2, we show the summary rules for each transaction type and how they contribute to the payout and payin lists.
@@
$\mathsf{Deposit}(\mathsf{type},\mathsf{amnt})$: allows a user to deposit an amount $\mathsf{amnt}$ of token with type $\mathsf{type}$ to be used for the next epoch.
@@
$\mathsf{Flash}(\mathsf{aux})$: Receive a flash loan request where $\mathsf{aux}$ contains all required inputs, then calculate the token amount the pool can provide and initiate the callback process (more details can be found in Section 4.2).
@@
Base smart contract $\mathsf{TokenBank}$. The mainchain part of the AMM is a base smart contract called $\mathsf{TokenBank}$. At an abstract level, as shown in Figure 4, this contract supports creating and managing token pools (i.e., tracking their balances and liquidity positions). It also provides the minimal interface needed to support users’ activities on the sidechain, which is mainly creating deposits containing the tokens they want to trade or provide as liquidity. This is needed since the sidechain does not receive or send actual tokens, it only tracks balance evolution. Hence, a user deposits the total amount of tokens they would need during an epoch before this epoch starts, and $\mathsf{TokenBank}$ handles the payouts and payins produced by the sidechain when the epoch ends.
@@
Sidechain management. The sidechain in $\mathsf{ammBoost}$ is managed in a similar way as in chainBoost. At the beginning of each epoch, a committee from the sidechain miners is elected, which runs a PBFT consensus protocol to agree on mining meta/summary blocks and issuing sync-transactions. That is, the committee leader proposes new blocks or sync-transactions, and collect votes from the committee members. Once a vote majority is reached, the new block is added to the sidechain or the sync-transaction is sent to the mainchain. In $\mathsf{ammBoost}$, a sync-transaction is basically a call to the function $\mathsf{Sync}$ in $\mathsf{TokenBank}$ shown in Figure 4.
$\mathsf{ammBoost}$ differs from chainBoost in the aspect that the sidechain has its own miner population. In other words, mainchain miners are not responsible for managing the sidechain, and even may not know that a sidechain exists in the first place. Executing the $\mathsf{TokenBank}$ contract is like executing any other contract deployed at the application layer of the mainchain. As such, sidechain miners must possess some mining power to establish Sybil-resistant identities to be used in the committee election process. Any secure PBFT protocol in which election is based on the mining power can be used here, e.g., the proof-of-stake based protocol in [48].
4.2 System Operation—Transaction Processing

### v4

@@ II Background and Challenges
We provide an overview of AMMs and the chainBoost framework that we use in the design of $\mathsf{ammBoost}$.
@@
Automated market makers (AMMs). AMMs build platforms for token trading powered by the users themselves. This is done by establishing liquidity pools such that a pool trades a pair of tokens, say tokens $A$ and $B$. Users are divided into: clients—token sellers and buyers, and liquidity providers (LPs). Providing liquidity comes from the sellers since buying token $A$ requires paying the price in token $B$ (and vice versa), and from LPs who add tokens to the pool and collect fees in return. Constant function market makers (CFMM) is a popular choice for computing the trading price in AMMs. This formula keeps the ratio of token reserves, and consequently prices, in the pool as balanced as possible to reduce price slippage.
@@
At a basic level, an AMM supports several transaction types: for trading, there are (exact input and exact output) swaps, and for liquidity management, there are mints, burns and collects that allow LPs to submit liquidity positions, collect their fees, and withdraw these positions, respectively. A liquidity position is a data structure in the AMM smart contract containing a position ID, the ID (e.g., a public key) of the owner, the amount of liquidity tokens the position owner provided, and the total amount of fees accrued so far. Collects are simple currency transfer transactions in which a position owner requests to withdraw (all or part of) the fee balance in their position. While burns allow a position owner to withdraw all liquidity balance and fees of their position and deleting the position. AMMs may provide additional services, such as flash loans [70] allowing clients to take advantage of arbitrage opportunities across different platforms. Furthermore, more sophisticated liquidity approaches are being adopted, e.g., concentrated liquidity [71] that enables defining a price range over which liquidity will be applied to address issues related to inefficient use of provided funds.
@@
The functionality of an AMM is commonly implemented as a set of smart contracts on top of a smart contract-enabled blockchain, where Ethereum is the dominant choice so far. These contracts create and manage the liquidity pools, and provide the API needed to interact with the AMM. AMMs on public blockchains have several financial and security issues [72, 73, 74], e.g., front-running attacks, sandwich attacks, miner/maximal extractable value, etc. Understanding and solving these issues are active research areas. We do not discuss these issues further since they are not the focus of this work; we target the storage cost and throughput of AMMs.
@@
The chainBoost framework. chainBoost [67] is a sidechain-based solution that reduces blockchain size and confirmation delays, and boost transaction throughput. It introduces a new sidechain architecture that has a mutual-dependence relation with the mainchain, which allows sharing the workload between the two chains, and supports pruning stale records. In particular, all service-related transactions that can be summarized go to the sidechain, while the rest stay on the mainchain. The sidechain works in parallel to the mainchain, and operates in epochs and rounds (an epoch is $\omega$ consecutive rounds and a round is the period during which a new block is mined). At the end of each epoch, the mainchain is synced with summaries of the workload processed by the sidechain in that epoch.
@@
The sidechain is managed by the mainchain miners, where for each epoch, a committee is elected to process the sidechain traffic. The rest of the mainchain miners, who are not on the committee, do not process the sidechain traffic, thus reducing their load. To speedup agreement, chainBoost employs a secure practical Byzantine fault tolerance (PBFT)-based consensus (similar to those in [75, 76]) for the sidechain.
@@
The sidechain is composed of two types of blocks: temporary meta-blocks and permanent summary-blocks. For each sidechain round, the committee mines a meta-block containing the transactions they processed, so that once a transaction appears in a meta-block it is considered final. In the last round of the epoch, this committee mines a summary-block summarizing all state changes induced by the meta-blocks within that epoch. After that, it issues a sync-transaction containing the summarized state changes, which the mainchain miners use to update the relevant state variables on the mainchain. Once the sync-transaction is confirmed on the mainchain, all meta-blocks used to produce the respective summary-block are discarded. This significantly reduces the sidechain size, and subsequently, the mainchain size. At the same time, having permanent summary-blocks enables anyone to verify the source of the state changes recorded on the mainchain.
Challenges of using dependent-sidechains for AMMs. AMMs’ setting is different from the one in [67]: First, in chainBoost the two chains are mutually-dependent and aware of each other. In $\mathsf{ammBoost}$, mainchain miners belong to a smart contract-enabled blockchain, and the AMM is simply an application deployed on that blockchain. Modifying the mainchain protocol to have its miners operate the sidechain would lead to a hard fork. To avoid that, the sidechain needs its own miners to run its protocol (indeed a miner can operate on both chains by running two protocols—one for each chain).
Second, in $\mathsf{ammBoost}$, actual tokens reside on the mainchain, while trading activities happen on the sidechain. However, in AMMs, a user who want to place a trade or add liquidity, must prove that they own actual tokens covering the activity by transferring the tokens to the AMM contract before the transaction is processed. Thus, a mechanism is needed to provide guarantees that users own actual tokens to backup transaction they issue on the sidechain. At the same time, transaction execution must be based off the pool balances from the mainchain, and a mechanism is needed to handle actual token payouts on the mainchain resulting from trading activities that took a place on the sidechain.
Second, the mainchain miners in $\mathsf{ammBoost}$ are not required to track the sidechain, and even they may not be aware of its existence. Nonetheless, $\mathsf{TokenBank}$ must only accept sync-transactions coming from the rightful sidechain epoch committees. Thus, $\mathsf{ammBoost}$ must implement an authentication mechanism to ensure that only legit syncs are accepted. Devising techniques to address these issues resemble the core novelty of $\mathsf{ammBoost}$ system design.
III Preliminaries
@@
Notation. We use $\lambda$ to denote the security parameter, $\mathtt{pp}$ to denote the system public parameters, $\mathcal{L}$ to denote a ledger (or blockchain), $\mathcal{L}_{\mathsf{mc}}$ to denote the mainchain ledger, and $\mathcal{L}_{\mathsf{sc}}$ to denote the sidechain ledger. $\mathcal{L}_{\mathsf{mc}}$ is the smart contract-enabled blockchain on top of which the AMM base smart contract is deployed, while $\mathcal{L}_{\mathsf{sc}}$ is the sidechain of the AMM ecosystem. Each party maintains a secret key $\mathsf{sk}$ and a public key $\mathsf{pk}$. Lastly, we use $\mathsf{PPT}$ for probabilistic polynomial time.
@@
System model. $\mathsf{ammBoost}$ involves a base smart contract representing the AMM on the mainchain, and a sidechain that processes most of the AMM workload. Anyone can join/leave the AMM at anytime, and these parties are known using their public keys. Participants are three types: clients $\mathcal{C}$ who only use the AMM trading services, liquidity providers $\mathcal{LP}$ who provide liquidity for the pools operated by the AMM, and miners $\mathcal{M}$ who maintain the AMM sidechain. We do not place any restrictions on the mainchain beyond being a secure smart contract-enabled blockchain. $\mathsf{ammBoost}$ operates in rounds and epochs (as defined earlier). The sidechain runs a secure PBFT-based consensus, e.g., [75, 76], in which a committee is elected for each epoch to run the agreement (we provide a background on PBFT and committee election in Appendix A). [^2] To simplify the presentation, we adopt a leader-based PBFT in which a leader proposes a block for the committee to agree on (as in [75]). Nonetheless, voting-based PBFT (as in [76]) can be used. In any epoch, the committee mines new blocks: temporary meta-blocks that record transactions, and permanent summary-blocks that summarize meta-blocks mined in an epoch. The committee also issues sync-transactions to sync the base AMM smart contract deployed on the mainchain. Accordingly, the $\mathsf{ammBoost}$ framework provides the following functionalities:
@@
Takes as input $\lambda$ and $\mathcal{L}_{\mathsf{mc}}$. It configures the public parameters $\mathtt{pp}$, and deploys a base AMM smart contract on $\mathcal{L}_{\mathsf{mc}}$. It outputs $\mathtt{pp}$ and the initial sidechain ledger $\mathcal{L}_{\mathsf{sc}}^{0}$ (the genesis block referencing the mainchain block containing the base contract).
@@
Takes as input $\mathtt{pp}$ and outputs the party’s initial local state $\mathsf{state}$, which contains a keypair $(\mathsf{sk},\mathsf{pk})$, and in case of miners, the current view of $\mathcal{L}_{\mathsf{sc}}$.
@@
  • $\mathsf{tx}_{\mathsf{collect}}$: Allows an LP to collect their fees.
  • $\mathsf{tx}_{\mathsf{burn}}$: Allows an LP to withdraw their liquidity.
@@
Takes as input the sidechain state $\mathcal{L}_{\mathsf{sc}}$, a new block $\mathsf{B}$ with type $\mathsf{btype}=\mathsf{meta}$ or $\mathsf{btype}=\mathsf{summary}$. It outputs 1 if $\mathsf{B}$ is valid based on the syntax/semantics of the block type, and 0 otherwise.
@@
Security model. We aim to develop a secure state growth control solution that preserves AMM correctness and security. $\mathsf{ammBoost}$ builds a sidechain, which is a blockchain, that interacts with the application layer of the mainchain through the base AMM smart contract. This sidechain must be a secure ledger that satisfies safety and liveness as defined in [77].
@@
A secure blockchain must record only valid transactions and blocks, thus its protocol is parameterized by predicates to verify transactions and blocks. For dApps, validity is governed by the code of their smart contracts, and miners ensure that the ledger state changes have been produced by a successful execution of this code. $\mathsf{ammBoost}$ reduces the AMM functionality deployed on the mainchain, and it processes most of the workload (following the same logic of the AMM) on the sidechain. Thus, in our security analysis, we show that $\mathsf{ammBoost}$ preserves the security and correct operation (i.e., safety and liveness) of the original AMM.
@@
Adversary model. We assume a secure mainchain as defined above. For the sidechain, we have honest miners who follow the protocol, and malicious miners controlled by the adversary who may behave arbitrarily. The adversary can deploy new miners or corrupt existing ones, without going above the threshold of faulty nodes required by the sidechain consensus protocol. As we use a PBFT-based consensus for the sidechain, it follows the threshold assumption for PBFT [75]; the adversary can corrupt up to $f$ parties, where a committee size is $23f+$ 2 and $22f+$ 2 votes are needed to reach an agreement. The adversary can see all messages and transactions sent in the system (since we deal with public permissionless blockchain systems) and can reorder these messages and delay them. As in [13, 75, 77, 78, 79, 80], we assume bounded-delay message delivery, so any sent message (or transaction) will be delivered within $\Delta$ time. We assume slowly-adaptive adversaries [81] that can corrupt miners only at the beginning of each epoch. Lastly, we deal with $\mathsf{PPT}$ adversaries.
IV System Design
@@
$\mathsf{ammBoost}$ changes the AMM deployment structure, as shown in Figure 1. The base smart contract is minimal; mainly tracking token balances of users and liquidity pools, while most transaction processing happen on the sidechain. Summaries of the sidechain processed traffic are used to sync the base smart contract. In this section, we present the design of $\mathsf{ammBoost}$ including system setup, architecture and operation, handling interruptions, and its security.
Figure 1: The $\mathsf{ammBoost}$ framework (Txs stands for transactions).
IV-A System Setup
@@
The setup phase, see Figure 2, mainly specifies the traffic split between the chains and the summary rules for the sidechain traffic, as well as the sidechain parameters such as the epoch length [^3] The epoch duration impacts syncing frequency. Short epochs mean more sync-transactions, which incurs more gas cost and may impact throughput—as they are processed by the mainchain, however users would receive their tokens faster compared to long epochs. We empirically study the impact of this parameter in Appendix E. and its consensus configuration (e.g., committee size). Also, this phase involves deploying the AMM base smart contract on the mainchain and creating the sidechain.
@@
$\mathsf{SystemSetup}(1^{\lambda},\mathcal{L}_{\mathsf{mc}})$: Takes as input $\lambda$ and the current mainchain state $\mathcal{L}_{\mathsf{mc}}$, and does the following:
@@ 1. Generate the sidechain configuration parameters:
  • All parameters of the sidechain consensus protocol.
@@
2. Deploy $\mathsf{TokenBank}$ on the mainchain.
@@
Traffic classification and summary rules. AMM transactions and operations are divided into two groups: pool management and trading. Creating and managing token pools, as well as dispensing tokens to clients and LPs, are done on the mainchain since these deal with actual tokens. Flashes are also handled by the mainchain since they require instant token dispensing rather than at the end of the epoch. The rest of the transactions, including swaps, mints, burns, and collects, are handled by the sidechain. In $\mathsf{ammBoost}$, the sidechain does not hold custody of tokens, it only tracks their balances based on the processed transactions. Thus, during each epoch, the sidechain produces two structures:
  • A payout list containing users’ public keys and the amount/type of tokens they should receive, which is simply the updated deposit balance at the end of an epoch.
The actual token dispensing and deduction happen at the end of an epoch when the sidechain summaries are received. The new state of the pool token balances on the mainchain will be computed based on these lists. Also, the payout encompasses refunding any leftover in the deposits to their owners. In Section IV-B, we show the summary rules for each transaction type and how they contribute to the payout and payin lists.
Base smart contract $\mathsf{TokenBank}$. The mainchain part of the AMM is a base smart contract called $\mathsf{TokenBank}$. At an abstract level, as shown in Figure 3, this contract supports creating and managing token pools (i.e., tracking their balances and liquidity positions). It also provides the minimal interface needed to support users’ activities on the sidechain, which is mainly creating deposits containing the tokens they want to trade or provide as liquidity. This is needed to guarantee that users own tokens since the sidechain does not hold actual tokens, it only tracks balance evolution. Hence, a user deposits the total amount of tokens they would need during an epoch before this epoch starts, and $\mathsf{TokenBank}$ handles the payouts and payins produced by the sidechain when the epoch ends.
@@
$\mathsf{Deposit}(\mathsf{type},\mathsf{amnt})$: allows a user to deposit an amount $\mathsf{amnt}$ of token with type $\mathsf{type}$ for the next epoch.
@@
$\mathsf{Flash}(\mathsf{aux})$: Receive a flash loan request where $\mathsf{aux}$ contains all required inputs, then calculate the token amount the pool can provide and initiate the callback process (more details can be found in Section IV-B).
@@
Sidechain management. Different from chainBoost, the sidechain in $\mathsf{ammBoost}$ is managed by its own miner population. Like any blockchain, anyone who can join as a miner given that they have mining power (e.g., a stake) to establish Sybil-resistant identities. By doing so, deploying $\mathsf{ammBoost}$ does not introduce core changes to the mainchain protocol (required if mainchain miners are to manage the sidechain), thus avoiding the consequences of having hard forks.
@@
At the beginning of each epoch, a committee from these miners is elected (e.g., using random sortition [76, 13]), which runs a PBFT consensus protocol to agree on mining meta/summary blocks and issuing sync-transactions. That is, the committee leader proposes new blocks or sync-transactions, and collect votes from the committee members. Once a vote majority is reached, the new block is added to the sidechain or the sync-transaction is sent to the mainchain. In $\mathsf{ammBoost}$, a sync-transaction is basically a call to the function $\mathsf{Sync}$ in $\mathsf{TokenBank}$ shown in Figure 3. Executing the $\mathsf{TokenBank}$ contract is like executing any other contract deployed at the application layer of the mainchain.
$\mathsf{ammBoost}$ does not place any restrictions on the consensus protocol of the underlying mainchain (beyond being a secure one), and the mainchain consensus protocol could be different from the sidechain one. Also, for $\mathsf{ammBoost}$’s sidechain, any secure PBFT-based consensus and any secure committee election algorithm can be used. Moreover, the committee size has a lower bound that guarantees the honesty threshold required by consensus (formal analysis can be found in [67]).
IV-B System Operation—Transaction Processing